OpenStudy (anonymous):

Find the perimeter of a rectangular object which has a length of feet and a width of feet

OpenStudy (anonymous):

length \[\sqrt{180}\] width \[\sqrt{125}\]

OpenStudy (anonymous):

you add the two lengths and the two widths

OpenStudy (anonymous):

so it would be \[\sqrt{180}\] + \[\sqrt{180}\] +\[\sqrt{125}\] + \[\sqrt{125}\] to simplify radicals you take find a perfect square that the number is divisble by, take the square root and then put it outside so \[\sqrt{125}\] = \[\sqrt{25times5}\] = 5 times\[\sqrt{5}\]